OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b (-4,9); 6x=5y+7

OpenStudy (anonymous):

Do you remember the point slope formula? That's a good start for you.

OpenStudy (anonymous):

6x=5y+c plug x=-4 and y=9 for finding c: 6(-4)=5(9)+c so, c=-69 so u get 6x=5y-69..

OpenStudy (anonymous):

k..i assume u know that equation of a line parallel to ax+by+c=0 is ax+by+k=0..where k and c are different constants..:)

OpenStudy (anonymous):

Remember parallel lines are parallel if they have the same slope (m). You have an equation for the first line. Perhaps you can rearrange the equation to find m. You also have a point. (x1, y1) = (-4, 9) According to the point slope formula: \[(y - y _{1}) = m(x - x _{1})\] You can plug in these values to find the line with a slope of m through a given point, (x1, y1)